4.1 Itata Limited has the choice of purchasing one of two machines viz. Machine L or Machine T. Both machines have a five-year life. The annual revenues from each machine are estimated at R2 000 000. Machine L is not expected to have a scrap value. Machine L costs R4 500 000. Its annual cash operating costs are estimated at R680 000. Machine T costs R4 500 000. Its annual cash operating costs are estimated at R700 000. The scrap value of this machine is estimated to be R200 000. Depreciation amounts to R900 000 per year for Machine L and R860 000 for Machine T. The cost of capital may be assumed to be 14%.

Required:

4.1.1 Calculate the Payback Period of Machine L (answer expressed in years, months and days). (3)

4.1.2 Use the Net Present Value technique to determine the machine that should be selected by Itata Limited. (7)

4.1.3 Calculate the Accounting Rate of Return of Machine L (answer expressed to two decimal places). (5) 3

4.2 The owner of Etihad Enterprises was approached by a local dealer in air conditioning units. The dealer proposed replacing the old cooling system of Etihad Enterprises with a modern, more efficient system. The cost of the new system was quoted at R300 000, but it would save R62 000 per year in electricity costs. The estimated life of the new system is 10 years, with no salvage value expected. All capital projects are required to earn at least the firm’s cost of capital, which is 12%.

Required:

Calculate the project’s Internal Rate of Return (IRR). (Round off your answer to two decimal places.) (10)
Answer #1

Calculation of Payback Period of Machine L-

Payback period = Investment/Cash Inflow

Cash inflow = Revenue - operating expenses=R2,000,000-680,000=R1,320,000

Payback period =R 4,500,000/R1320000=3.41 years

Payback period = 3years + (0.41×12)/1= 3yrs + 4.92month =3yrs +4months +( 0.92×365)/12

Payback period is 3years,4months and 27days.

Q4.1.1

Machine L payback period is 3yrs4months 27days

Q4.1.2

Machine T Payback period

Investment = cash outflow - scrap value

Investment = R4,500,000-R200,000=R4,300,000

Cashinflow = Revenues - Operating expenses=R 2,000,000-R700,000=R1,300,000

payback period of Machine T = R4,300,000/R1,300,000=3.31yrs

3yrs + 0.31×12months = 3yrs+ 3.72mnths = 3yrs + 3month +0.72×365/12

3year3months 22days.

Machine T Payback period is faster than machine L

Hence, MACHINE T should be selected.

Note:

[Depreciation is non cash expenses.It should be excluded from operating expenses]

Q4.1.3

Calculation of Accounting rate of return of Machine L

Accounting Rate of Returns = Net profit / investment

Annual profit = Revenue -(Operating expenses + Depreciation)

= R2000,000-(R680,000+R900,000)

=R2,000,000-R1,580,000=R420,000

ARR= R(420,000/R4,500,000)×100= 9.33%

___________________________________________________

​​​​​​Post separately Q4.2